# Repurposing Pre-trained Video Diffusion Models for Event-based Video Interpolation

### v1

@@ 4.3 Evaluation on Video Frame Interpolation
Figure 8: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Camera_Far, which involves large camera motion capturing nearby objects, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)
@@
Figure 11: The qualitative visualization of our experimental results on the BS-ERGB test sequences, with 3 skips between the start and end frames. All interpolated frames are presented. This sequence features fast motion of a hand and a rising, spinning ball, shows that our method is the only approach capable of maintaining the regular texture shapes on both the ball and hand across the interpolated frames in this challenging scenario. Please refer to the website for additional results on datasets and videos, the videos provide further and straightforward comparisons of reconstruction consistency.
5 Limitations
Despite the promising results, adapting pre-trained video diffusion models to the EVFI task presents several limitations. First, the pre-trained VAE encoder and decoder, used for conversion between pixel space and latent space, have limited representational power, often resulting in color shifts and loss of small details. Second, the adapted video diffusion model tends to smooth out certain regions, likely due to its difficulty in preserving fine textures within the downsampled latent space or the need for a prohibitively high upsampling factor to prevent smoothing.
These limitations stem from the current state of pre-trained video diffusion foundation models. However, as the field of video generation evolves rapidly, we anticipate that many of these issues could be mitigated or resolved with newer, more advanced pre-trained models in the near future.
6 Conclusion

### v2

@@ 14 Model Run Time, Memory, and Parameter Comparison
Figure 16: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Camera_Far, which involves large camera motion capturing distant objects, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)
Figure 17: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Camera_Close, which involves large camera motion capturing nearby objects, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)